OpenStudy (anonymous):

WILL REWARD Which cosine function has maximum of 4, a minimum of -4, and a period of 2pi/3? A. y=4 cos 3 theta B. y= 4 cos 2 theta/3 C. y=4 cos theta/3 D. y=4 cos 3 theta

OpenStudy (anonymous):

Since \(\cos\) ranges between -1 and 1, it's really easy to see that any \(4\cos\) will have a min and max of -4 and 4 respectively. A function \(\cos(B \times \theta)\) has period \(\frac{2\pi}{B}\). So if we want the period to be \(\frac{2\pi}{3}\), \(B=3\). Putting both conditions together, the function must be \(4\cos(\frac{\theta}{3})\), choice C.